# Infer-and-widen, or not?

### v2

@@ 4 Vignette #2: Maximal contrasts
Following Zrnic and Jordan (2023), we consider the model 3.1 and suppose we also have access to a fixed design matrix with normalized columns $\left\|X_{1}\right\|_{2}=\dots=\left\|X_{p}\right\|_{2}=1$. Our interest lies in the quantity $X_{\tilde{\mathcal{S}}(Y)}^{\top}\mu$ defined via the selection rule
$\tilde{\mathcal{S}}(Y):=\mathop{\mathrm{arg\,max}}_{j\in[p]}\left|X_{j}^{\top}Y\right|;$ (4.1)
Zrnic and Jordan (2023) refer to this as the “maximal contrast”. If $X=I$, then this setting is similar to Vignette #1, except with an absolute value in the selection rule.
To achieve a “stable” algorithm, Zrnic and Jordan (2023) consider the selection rule
$\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y):=\mathop{\mathrm{arg\,max}}_{j\in[p]}\left|X^{\top}_{j}Y+\zeta_{j}\right|,$ (4.2)
@@
Proposition 1 (An infer-and-widen interval based on (4.2), given by Zrnic and Jordan (2023)).
Under (3.1), consider the Laplacian selection rule $\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)$ defined in (4.2). For any $\eta>0$ and $\nu\in(0,1)$ such that $c\geq 2z_{1-\alpha\nu/(2p)}/\eta$, the infer-and-widen confidence interval
$\mathrm{CI}_{{{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}}(Y)}}^{\alpha(1-\nu)e^{-\eta}}(Y):=\left[X^{\top}_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}Y\pm\sigma z_{1-\alpha(1-\nu)e^{-\eta}/2}\right]$ (4.3)
has $1-\alpha$ unconditional selective coverage for $X^{\top}_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}\mu$ in the sense of 2.1.
We will take a randomized conditional selective inference approach (Tian and Taylor, 2018; Panigrahi et al., 2024) to construct an alternative confidence interval that makes use of the identical selection event.
@@
Consider the model in 3.1 and selection rule $\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)$ in 4.2. Define
$W$ $:=(I_{n}-X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}^{\top})Y,$
$\delta$ $:={\rm sign}\left(X^{\top}_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}Y+\zeta_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}\right),$
@@
where $\Phi_{\lambda,\sigma^{2},[v_{\min},v_{\max}]}(\cdot)$ is defined in Proposition 2, and $v_{\min}$ and $v_{\max}$ are defined in LABEL:sec:deferred. The confidence interval
$\mathrm{CI}_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}^{\alpha}\left(Y\right):=\left[a_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)},b_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}\right]$ (4.6)
@@
has $1-\alpha$ conditional selective coverage for $X_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}^{\top}\mu$ in the sense of 1.3 with $\mathcal{C}(Y)=\left(\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y),\zeta,W,\delta\right)$.
@@
Computing 4.6 requires numerically solving 4.4 for $a_{i}$ and 4.5 for $b_{i}$, where $i=\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)$.
Figure 2: Vignette #2 under (3.1) with $n=100$, $\sigma=1$, $\mu=X\phi$, and the selection rule in 4.2, where $X$ is multivariate normal with column correlation $0.5$, and $\phi$ is sparse with exponentially distributed non-zero elements. Results are averaged over (a) $250$ or (b, c) $1000$ simulated datasets. (a): We display the ratio of the average widths of infer-and-widen 4.3 and randomized conditional selective inference (RCSI) 4.6 intervals with $95\%$ coverage across the number of features $p$ and Laplacian noise variance $2c^{2}$. A width ratio that exceeds one implies that the infer-and-widen interval is wider. Black regions represent pairs of $n$ and $c$ for which the infer-and-widen approach requires such a small Gaussian quantile that R returns an interval of infinite width. (b, c): Fixing $p=100$, we display the average width of the “oracle” infer-and-widen interval and the RCSI interval 4.6 as a function of coverage. The 95% classical interval and the “attainable” 95% infer-and-widen interval 4.3 based on algorithmic stability are also displayed. The Laplacian noise variance is $0.33$ and $3$ in panels (b) and (c), respectively.
@@
We simulate data according to 3.1 with $n=100$ and $\sigma=1$. We set $\mu=X\phi$ where half of the elements of $\phi$ are zero and the other half are drawn from an exponential distribution with mean $50/7$. The rows of $X$ are independent draws from ${\bf{\rm N}}\,_{p}(0,\rho(1_{p}1_{p}^{\top}-I_{p})+I_{p})$ with $\rho=0.5$, and its columns are subsequently scaled so that $\left\|X_{1}\right\|_{2}=\dots=\left\|X_{p}\right\|_{2}=1$. Figure 2(a) displays the ratio of confidence interval widths across values of $p$ and the variance of the Laplacian noise $2c^{2}$ in 4.2.
The infer-and-widen interval (4.6) involves tuning parameters $\nu$ and $\eta$; to construct Figure 2(a), for a fixed value of the noise $c$ and number of features $p$ in (4.2), we select the $\nu$ and $\eta$ that minimize the confidence interval width, subject to the constraints in Proposition 1. In the figure, black indicates values of $c$ and $p$ that necessitate a tiny quantile of the normal distribution, yielding numerical issues in R and therefore an interval (4.3) of infinite width. The randomized conditional selective inference intervals are narrower than the infer-and-widen intervals across most values of $p$ and $c$ considered.
@@
Once again, we consider the “oracle” infer-and-widen interval: this is the narrowest possible interval centered at $X_{\tilde{\mathcal{S}}_{c}^{\mathrm{Laplace}}(Y)}^{\top}Y$ that achieves a given coverage. In Figures 2(b, c), we numerically compare the “oracle” infer-and-widen method to the randomized conditional selective inference interval (4.6) under the selection rule (4.2) with variances $0.332c^{2}=$ 0.33 and $3$, respectively. In this setting, the “oracle” infer-and-widen interval is narrower than the randomized conditional selective inference interval in (4.6), thereby indicating that in principle, an infer-and-widen interval can outperform the randomized conditional selective inference interval (4.6) in this setting. However, to the best of our knowledge, no methods to construct such a narrow infer-and-widen interval are currently available.
@@ 5 Vignette #3: Inference after the lasso
We now consider inference on the set of coefficients selected by the lasso in a linear model (Tibshirani, 1996). We again consider the model 3.1, alongside a fixed design matrix $X\in\mathbb{R}^{n\times p}$ with unit-normalized columns, i.e., $\left\|X_{1}\right\|_{2}=\dots=\left\|X_{p}\right\|_{2}=1$. Our new selection rule is the magnitude-ordered set of features selected by the lasso:
$\bar{\mathcal{S}}(Y)=$$\bigg{\{}j_{1},\dots,j_{l}:\left|\hat{\beta}_{j_{1}}^{\lambda}\right|>\dots>\left|\hat{\beta}_{j_{l}}^{\lambda}\right|>0,\;$ (5.1)
$\quad\underset{k\notin\{j_{1},\ldots,j_{l}\}}{\max}|\hat{\beta}_{k}^{\lambda}|=0\bigg{\}},$
@@
$\hat{\beta}^{\lambda}:=\mathop{\mathrm{arg\,min}}_{\beta}\left\{\left\|Y-X\beta\right\|^{2}_{2}+\lambda\|\beta\|_{1}\right\}.$
We then consider a least squares model fit to the features in $\bar{\mathcal{S}}(Y)$. Our target of inference is the coefficient of the feature with largest lasso magnitude, i.e., $\beta_{{\bar{\mathcal{S}}(Y)},1}$, where
$\beta_{\bar{\mathcal{S}}(Y)}:=A_{\bar{\mathcal{S}}(Y)}\mathbb{E}\left[Y\right]$
and
$A_{J}:=\left(X_{J}^{\top}X_{J}\right)^{-1}X_{J}^{\top}$
for any subset of features $J\subseteq[p]$.
The lasso is not stable in the sense of (Zrnic and Jordan, 2023). Therefore, Zrnic and Jordan (2023) consider a variant of the lasso that is far enough from the original lasso proposal that we do not investigate it here. Instead, in what follows, we construct an interval based on Gaussian data thinning (Neufeld et al., 2024; Dharamshi et al., 2024) that yields conditional guarantees in the sense of 1.3, and then consider the “oracle” infer-and-widen interval that makes use of the identical selection event.
Towards this end, for some $c>0$, we define
$\bar{\mathcal{S}}_{c}^{\mathrm{Gaussian}}(Y)=\bar{\mathcal{S}}(Y+c\zeta),\quad\zeta_{i}\stackrel{{\scriptstyle{i.i.d.}}}{{\sim}}{\bf{\rm N}}\,(0,\sigma^{2}).$ (5.2)
Proposition 1.
Under (3.1), consider the selection rule $\bar{\mathcal{S}}_{c}^{\mathrm{Gaussian}}(\cdot)$ defined in (5.2). The data thinning confidence interval
$\mathrm{CI}_{\bar{\mathcal{S}}_{c}^{\mathrm{Gaussian}}(Y)}^{\alpha}(Y):=\left[\hat{\beta}_{\bar{\mathcal{S}}_{c}^{\mathrm{Gaussian}}(Y),1}\pm z_{1-\alpha/2}\left(\Sigma_{1,1}\right)^{1/2}\right]$ (5.3)
has $1-\alpha$ conditional selective coverage for $\beta_{\bar{\mathcal{S}}_{c}^{\mathrm{Gaussian}}(Y),1}$ in the sense of 1.3 with $\mathcal{C}(Y):=\left(Y+c\zeta\right),$ where
$\hat{\beta}_{\bar{\mathcal{S}}_{c}^{\mathrm{Gaussian}}(Y)}:=A_{\bar{\mathcal{S}}_{c}^{\mathrm{Gaussian}}(Y)}\left(Y-\tfrac{1}{c}\zeta\right)$
and
$\Sigma:=\sigma^{2}\left(1+\tfrac{1}{c^{2}}\right)\left(X_{\bar{\mathcal{S}}_{c}^{\mathrm{Gaussian}}(Y)}^{\top}X_{\bar{\mathcal{S}}_{c}^{\mathrm{Gaussian}}(Y)}\right)^{-1}.$
We compare the widths of the data thinning (DT) (5.3) and “oracle” infer-and-widen intervals — with identical selection events — in Figure 3. We simulate data according to 3.1 with $n=100$ and $\sigma=1$. We set $\mu=X\phi$, where half of the elements of $\phi$ are zero and the other half are either zero or drawn from exponential distributions with means $50/7$ or $5$ in panels (a), (b), and (c) of Figure 3. The rows of $X$ are independent draws from ${\bf{\rm N}}\,_{p}(0,\rho(1_{p}1_{p}^{\top}-I_{p})+I_{p})$ with $\rho=0.5$, and the columns of $X$ are scaled so that $\left\|X_{1}\right\|_{2}=\dots=\left\|X_{p}\right\|_{2}=1$. The penalty $\lambda$ is adaptively chosen using 3-fold cross-validation. Figure 3 displays the ratio of the average confidence interval widths as we vary $p$ and the variance of the Gaussian noise $c^{2}$ in (5.2).
We see from Figure 3 that the relative widths of the two intervals depend on the variance of the noise in (5.2) and distribution of $\phi$. Of course, the “oracle” infer-and-widen interval cannot be computed in practice: any real-world infer-and-widen interval that achieves nominal coverage would necessarily be at least as wide as the oracle interval, if not far wider. Thus, while it may be possible to develop an infer-and-widen interval that might outperform the simple data thinning interval in 5.3 in specific scenarios, the potential upside of any such approach is extremely limited.
Figure 3: Vignette #3 under (3.1) with $n=100$, $\sigma=1$, $\mu=X\phi$, and the selection rule in 5.2. The columns of $X$ are independent normal draws with column correlation $0.5$. Results are averaged over $1000$ simulated datasets. We consider inference on a lasso-selected coefficient in a least squares model, varying the mean of $\phi$ across $0,5,50/7$ in panels (a), (b), and (c), respectively. We display the ratio of average widths of the 95% “oracle” infer-and-widen interval and the 95% data thinning (DT) interval (5.3), with identical selection events and level $\alpha=0.05$; a positive log ratio indicates that the data thinning interval is narrower and hence preferable. In many settings, even the unattainable “oracle” infer-and-widen interval is wider than the attainable data thinning interval.

### v3

@@ 4 Vignette #2: Maximal contrasts
Following Zrnic and Jordan [28], we consider the model 3.1 and suppose we also have access to a fixed design matrix with normalized columns $\left\|X_{1}\right\|_{2}=\dots=\left\|X_{p}\right\|_{2}=1$. Our interest lies in the quantity $X_{\mathcal{S}(Y)}^{\top}\mu$ defined via the selection rule
$\mathcal{S}(Y):=\mathop{\mathrm{arg\,max}}_{j\in[p]}\left|X_{j}^{\top}Y\right|;$ (4.1)
[28] refer to this as the “maximal contrast”. If $X=I$, then this setting is similar to Vignette #1, except with an absolute value in the selection rule.
To achieve a “stable” algorithm, [28] consider the selection rule
$\mathcal{S}_{c}^{\mathrm{Laplace}}(Y):=\mathop{\mathrm{arg\,max}}_{j\in[p]}\left|X^{\top}_{j}Y+\zeta_{j}\right|,$ (4.2)
@@
Proposition 3 (An infer-and-widen interval based on (4.2), given by Zrnic and Jordan [28]).
Under (3.1), consider the Laplacian selection rule $\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)$ defined in (4.2). For any $\eta>0$ and $\nu\in(0,1)$ such that $c\geq 2z_{1-\alpha\nu/(2p)}/\eta$, the infer-and-widen confidence interval
$\mathrm{CI}_{{{\mathcal{S}_{c}^{\mathrm{Laplace}}}(Y)}}^{\alpha(1-\nu)e^{-\eta}}(Y):=\left[X^{\top}_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}Y\pm\sigma z_{1-\alpha(1-\nu)e^{-\eta}/2}\right]$ (4.3)
has $1-\alpha$ unconditional coverage for $X^{\top}_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}\mu$ in the sense of 2.1.
We will take a randomized conditional inference approach [24, 20] to derive an alternative confidence interval that makes use of the identical selection event.
@@
Consider the model in 3.1 and selection rule $\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)$ in 4.2. Define
$W$ $:=(I_{n}-X_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}X_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}^{\top})Y,$
$\delta$ $:={\rm sign}\left(X^{\top}_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}Y+\zeta_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}\right),$
@@
where $\Phi_{\lambda,\sigma^{2},[v_{\min},v_{\max}]}(\cdot)$ is defined in Proposition 2, and $v_{\min}$ and $v_{\max}$ are defined in Appendix A. The confidence interval
$\mathrm{CI}_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}^{\alpha}\left(Y\right):=\left[a_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)},b_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}\right]$ (4.6)
@@
has $1-\alpha$ conditional coverage for $X_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}^{\top}\mu$ in the sense of 1.3 with $\mathcal{C}(Y)=\left(\mathcal{S}_{c}^{\mathrm{Laplace}}(Y),\zeta,W,\delta\right)$.
@@
Computing 4.6 requires numerically solving 4.4 for $a_{i}$ and 4.5 for $b_{i}$, where $i=\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)$.
@@
We simulate data according to 3.1 with $n=100$ and $\sigma=1$. We set $\mu=X\phi$ where half of the elements of $\phi$ are zero and the other half are drawn from an exponential distribution with mean $50/7$. The rows of $X$ are independent draws from ${\bf{\rm N}}\,_{p}(0,\rho(1_{p}1_{p}^{\top}-I_{p})+I_{p})$ with $\rho=0.5$, and its columns are subsequently scaled so that $\left\|X_{1}\right\|_{2}=\dots=\left\|X_{p}\right\|_{2}=1$. Figure 2(a) displays the ratio of confidence interval widths across values of $p$ and the variance of the Laplacian noise $2c^{2}$ in 4.2. The infer-and-widen interval (4.6) involves tuning parameters $\nu$ and $\eta$; to construct Figure 2(a), for a fixed value of the noise $c$ and number of features $p$ in (4.2), we select the $\nu$ and $\eta$ that minimize the confidence interval width, subject to the constraints in Proposition 3. The randomized conditional selective inference intervals are narrower than the infer-and-widen intervals across most values of $p$ and $c$ considered.
@@
Once again, we consider the “oracle” infer-and-widen interval: this is the narrowest possible interval centered at $X_{\mathcal{S}_{c}^{\mathrm{Laplace}}(Y)}^{\top}Y$ that achieves a given coverage. In Figures 2(b, c), we numerically compare the “oracle” infer-and-widen method to the randomized conditional selective inference interval (4.6) under the selection rule (4.2) with variances $2c^{2}=0.33$ and $3$, respectively. In this setting, the “oracle” infer-and-widen interval is narrower than the randomized conditional selective inference interval in (4.6), thereby indicating that in principle, an infer-and-widen interval can outperform the randomized conditional selective inference interval (4.6) in this setting. However, to the best of our knowledge, no methods to construct such a narrower infer-and-widen interval are currently available.
Figure 2: Vignette #2 under (3.1) with $n=100$, $\sigma=1$, $\mu=X\phi$, and the selection rule in 4.2, where $X$ is multivariate normal with column correlation $0.5$, and $\phi$ is sparse with exponentially distributed non-zero elements. Results are averaged over (a) $250$ or (b, c) $1000$ simulated datasets. (a): We display the ratio of the average widths of infer-and-widen (I&W) 4.3 and randomized conditional selective inference (RCSI) 4.6 intervals with $95\%$ coverage across the number of features $p$ and Laplacian noise variance $2c^{2}$. (b, c): Fixing $p=100$, we display the average width of the “oracle” infer-and-widen interval and the RCSI interval 4.6 as a function of coverage. The 95% classical interval and the “attainable” 95% infer-and-widen interval 4.3 based on algorithmic stability are also displayed. The Laplacian noise variance is low and high in panels (b) and (c), respectively.
@@ 5 Vignette #3: Inference after the lasso
We now consider inference on the set of coefficients selected by the lasso in a linear model. We again consider the model 3.1, alongside a fixed design matrix $X\in\mathbb{R}^{n\times p}$. Our selection rule is the set of features selected by the lasso estimator $\hat{\beta}^{(\lambda)}$ [25]:
$\mathcal{S}(Y):={\bf{\rm supp}}\;\left(\hat{\beta}^{(\lambda)}\right),$ (5.1)
@@
$\hat{\beta}^{(\lambda)}:=\mathop{\mathrm{arg\,min}}_{\beta}\left\{\tfrac{1}{n}\left\|Y-X\beta\right\|^{2}_{2}+\lambda\|\beta\|_{1}\right\}.$
This is a “non-randomized” selection event, and we defer to Section E of the supplement for inference after a related randomized selection event.
Our goal is valid inference on the least squares coefficients in the model selected by $\mathcal{S}(Y)$:
$\beta_{\mathcal{S}(Y)}:=\left(X_{\mathcal{S}(Y)}^{\top}X_{\mathcal{S}(Y)}\right)^{-1}X_{\mathcal{S}(Y)}^{\top}\mathbb{E}\left[Y\right],$
where $X_{\mathcal{S}(Y)}$ is the subset of columns of $X$ indexed by $\mathcal{S}(Y)$. Since $\mathcal{S}(Y)$ is a data-adaptive model, the naive point estimator
$\hat{\beta}_{\mathcal{S}(Y)}:=\left(X_{\mathcal{S}(Y)}^{\top}X_{\mathcal{S}(Y)}\right)^{-1}X_{\mathcal{S}(Y)}^{\top}Y$
is no longer guaranteed to be unbiased for $\beta_{\mathcal{S}(Y)}$.
Obtaining valid inference on $\beta_{\mathcal{S}(Y)}$ in the sense of 1.2 is a classic problem in the field of selective inference, and there are two main categories of methods which do so. The first category guarantees unconditional coverage, 1.2, via the stronger goal of simultaneous coverage. Since the Bonferroni correction [8] is valid but would require an overly conservative adjustment for all $O(2^{p})$ possible models, Berk et al. [3] derive the narrower PoSI simultaneous correction. Zrnic and Fithian [26] subsequently developed the locally simultaneous inference approach which focuses the simultaneous correction on only a subset of likely models, thereby often improving power; these two approaches yield infer-and-widen intervals centered at $\hat{\beta}_{\mathcal{S}(Y)}$ and with widths proportional to the diagonal elements of $\left(X_{\mathcal{S}(Y)}^{\top}X_{\mathcal{S}(Y)}\right)^{-1/2}$ with a shared constant of proportionality. The second category of methods approach unconditional coverage via the stronger goal of conditional coverage, 1.3. The foundational work of Lee et al. [16] characterized the selection event $\mathcal{S}(Y)$, and used this to derive the conditional distribution $\hat{\beta}_{\mathcal{S}(Y)}\mid\mathcal{C}(Y)$ for inference on $\beta_{\mathcal{S}(Y)}$, where the sigma algebra of $\mathcal{C}(Y)$ is a superset of that of ${\mathcal{S}}(Y)$. In order to improve upon the power of this method, McCloskey [18] developed the hybrid method, which uses the conditional confidence interval in cases where there is strong signal, and otherwise resorts to unconditional intervals; thus the hybrid method provides only unconditional and not conditional coverage.
As before, we compare the widths of these methods’ confidence intervals, as well as the “oracle” infer-and-widen interval: the narrowest possible interval centered at $\hat{\beta}_{\mathcal{S}(Y)}$ and with width proportional to the diagonal elements of $\left(X_{\mathcal{S}(Y)}^{\top}X_{\mathcal{S}(Y)}\right)^{-1/2}$ with a shared constant of proportionality that attains coverage at a given level. We simulate data according to 3.1 with $n=100$, $p=10$, $\sigma=1$, and fixed penalty $\lambda$. We set $\mu=X\beta$, where $\beta_{1}=\dots=\beta_{5}$ and $\beta_{6}=\dots=\beta_{10}=0$. The rows of $X$ are independent draws from ${\bf{\rm N}}\,_{p}(0,\rho(1_{p}1_{p}^{\top}-I_{p})+I_{p})$ with correlation $\rho$. Figure 3(a) displays the ratio of average $95\%$ confidence interval widths of the locally simultaneous inference (LSI) [26] and hybrid [18] methods across values of $\rho$ and $\beta_{1}/\lambda$. A ratio greater than one indicates that the LSI method has a wider average interval width. In Figure 3(b, c), we fix $\rho=0.9$, and set $\beta_{1}/\lambda=0$ in panel (b) and $\beta_{1}/\lambda=1$ in panel (b). We numerically compare the “oracle” interval to the conditional [16] and hybrid [18] intervals. For reference, we also display the average empirical widths and coverages of the $95\%$ intervals from the classic, simultaneous PoSI (SI) [3], and LSI [26] approaches. In this setting, the “oracle” infer-and-widen interval is the narrowest interval, thereby indicating that in principle, an infer-and-widen interval can outperform the hybrid interval. However, the hybrid method outperforms both existing infer-and-widen methods.
Figure 3: Vignette #3 under (3.1) with $n=100$, $p=10$, $\sigma=1$, $\mu=X\beta$, and the selection rule in 5.1, where the rows of $X$ are equi-correlated normally distributed, and $\beta$ has five nonzero elements. Results are averaged over $100$ simulated datasets. (a): We display the ratio of the average widths of $95\%$ confidence intervals from the LSI infer-and-widen method [26] and the hybrid method [18]. A width ratio that exceeds one implies that the infer-and-widen interval is wider. (b, c): Fixing $\rho=0.9$, we compare the average width of the “oracle” infer-and-widen (I&W) interval to the conditional [16] and hybrid [18] intervals. For reference, we also display the average empirical widths and coverages of the $95\%$ intervals from the classic, simultaneous PoSI (SI) [3], and LSI [26] approaches. Panels (b) and (c) contain zero and nonzero signal, respectively.